# Mechanism and Stability of Li-Dynamics in Amorphous Li-Ti-P-S Based Mixed Ionic-Electronic Conductor

### v1

@@ Abstract
Mixed ionic-electronic conductor (MIEC) exhibit both high ionic and electronic conductivity to improve the battery performance. In this work, we investigate the mechanism and stability of transport channels in recently developed our MIEC material, amorphous Ti-doped lithium phosphorus sulfide (LPS), using molecular dynamics (MD) simulations with a 99% accurate machine-learning force field (MLFF) trained on ab-initio MD data. The achieved MLFF helps efficient large-scale MD simulations on LPS with three Ti concentrations (10%, 20%, and 30%) and six temperatures (25oC to 225oC) to calculate ionic conductivity, activation energy, Li-ion transport mechanism, and configurational entropy. Results show that ionic conductivities and activation energies are consistent with our recent experimental values. Moreover, Li-ion transport occurs via free-volume diffusion facilitated by the formation of disordered Li-S polyhedra. Enhanced stability of transport channels at 10% and 20% Ti doping, compared to 0% and 30%, is observed through the analysis of the vibrational and configurational entropy of these disordered Li-S polyhedra. Overall, this study highlights the utility of MLFF-based large-scale MD simulations in explaining the transport mechanism and its stability of Li-ion in Ti doped LPS electrolyte with significant computational efficiency.
@@ 1 Introduction
A MIEC in crystalline or amorphous state represent a promising class of materials for next-generation energy storage devices, including lithium-ion and lithium-sulfur (Li-S) batteries for more than one decade[7, 9, 10]. These materials exhibit both ionic and electronic conductivity, enabling them to play a dual role in facilitating lithium-ion transport while simultaneously providing pathways for electronic charge transfer. Unlike traditional solid electrolytes, which are purely ionic conductors, MIEC improve overall cell efficiency by addressing critical challenges such as sluggish redox kinetics, insulating discharge products, and interfacial instabilities[11]. Especially, the amorphous structure further enhances their ionic conductivity by reducing crystallographic barriers to ion migration, making them highly attractive for solid-state applications[12].
@@
Our approach in this work involves the integration of DFT based $ab-initio$ molecular dynamics (AIMD), machine learning force filed based on deep neural networks, and classical molecular dynamics. This integrated approach is collectively referred to as deep learning molecular dynamics (DLMD) simulations and its force field with potential energy surface called deep learning potential (DLP) [17, 18] to investigate the structural and Li-ion dynamics in Li-Ti-P-S electrolyte. With this, we organize the manuscript as follows: Section 2 describes detailed simulation techniques, as illustrated in Fig. 1. Section 3 is composed of results and discussions of mean square displacement (MSD), diffusion coefficient, ionic conductivity, activation energy, and transport mechanism and stability of Li-ion ions. Finally, we conclude our results and discussions in Section 4.
Figure 1: (a) Schematic diagram of the protocol of DLMD development (see Section 2.3), where R, F, E, and DL represents atomic coordinates, forces, energies, and deep learning, respectively. (b-e) Ball and stick models of amorphous LPS(b), LPS:Ti10%(c), LPS:Ti20%(d), and LPS:Ti30%(e), respectively. (f-j) Accuracy of trained force field model with DFT data versus predicted data such as energy(f) and forces along Fx, Fy, and Fz of Li(g), P(h), S(i), and Ti(j).
@@ 2.3 DLMD simulation
The training process consisted of $2\times 10^{4}$ steps, utilizing mean squared error for both energies and forces as the loss function at each training step. The optimization was facilitated by the Adam optimizer, initiating with a learning rate of $10^{-3}$ and concluding at $10^{-8}$, with a decay parameter set to 5000. With these specified parameters, a dataset comprising more than 2 million training samples and 20000 test samples from diverse trajectories were employed to construct the DLP model. The calculated loss function parameters such as mean absolute error and root mean square error during validation of developed DLP model is tabulated in Table-1. Accuracy of the predicted data of force and energy is more than 99 % and data points of predicted versus trained are shown in Fig. 1f to Fig. 1j.
@@ 3.1 Structure optimization
Using the developed DLP, DLMD simulations were performed in the isothermal-isobaric (NPT) ensemble to control the temperature and pressure of the system and optimize a large structure consisting of 12,000 atoms, designed from the output of AIMD. The simulation process began with an initial phase at a low temperature of 1 K for 2 ps to stabilize the initial configuration. Subsequently, the system was gradually heated from 1 K to 1000 K over another 2 ps, allowing a smooth transition to a higher thermal state. After reaching 1000 K, the system was equilibrated at this elevated temperature for 3 ps to ensure thermal and pressure stability. A controlled cooling phase was then implemented, reducing the temperature from 1000 K to 300 K over 3 ps. Finally, the system was equilibrated at 300 K for 50 ps to simulate standard room-temperature conditions. The temperature, potential energy, pressure, and volume profiles with respect to time for LPS:Ti00%,10%,20%,30% systems are plotted in Fig. 2 (a-d) and there optimized structures were shown in Fig. 2(e-h). All there simulations were done using LAMMPS simulation package [28, 29, 30] coupled with the DeepMD plugin [31].
With the optimized large structure, DLMD simulations were conducted in the NVT ensemble to investigate the Li-dynamic properties. Simulations were performed at eight different temperatures—300, 320, 340, 360, 380, 400, 450, and 500 K (25oC to 225oC)—controlled using the Nosé-Hoover thermostat. The temperature damping parameter was set to 100 fs, and a uniform integration time step of 1 fs was employed for all simulations, extending over a total duration of 3.0 ns. The small and large cells from the AIMD and DLMD simulations were visualized using VESTA software[32], as shown in Fig. 1 (b-e) and Fig. 2 (e-h), respectively. Additionally, the trajectories of the AIMD and DLMD simulations were analyzed and visualized using OVITO software.[33]
@@ 3.3 Ionic Conductivity and Activation Energy
Figure 5: Ball and stick model for moving Li-ion(Red, Li*) with respect to time step of LPS:Ti00%(a), LPS:Ti10%(b), LPS:Ti20%(c), and LPS:Ti30%(d). Here, cubic boundaries of the ball and stick models are sudo-boundaries that are only to show the movement of one particular Li-ion in a large-cell. The Li-ion movement in 2D plane ($xy$ and $xz$ in Å)with respect to time steps are plotted in (e-l).
@@ 3.4 Li-ion Transport Mechanism
Furthermore, the free-volume diffusion process in LPS involves the formation of disordered Li-Sn ($n$ = 1 to 6) polyhedra within the electrolyte matrix. The disordered nature of these polyhedra reduces structural constraints that results in variations in the migration speed of Li-ions and affecting the stability of Li-ion migration channels. Therefore, a quantitative analysis of Li-S polyhedra in Ti-doped LPS is crucial for understanding the stability and effectiveness of Li-ion transport channels.
@@ 3.5.2 Configurational Entropy
where $S_{\text{trans}}$ is the translational entropy, $S_{\text{vib}}$ is the vibrational entropy, $S_{\text{elec}}$ is the electronic entropy, and $S_{\text{config}}$ is the configurational entropy.
Translational and electronic entropies are negligibly small. The electronic and configurational entropy in Equation 8 was derived from the probabilities of different coordination environments using the Boltzmann entropy formula:
@@
The configurational entropy $S_{\text{config}}$ defined degree of disordered Li-S polyhedrals in LPS plays a significant role in determining the stability and dynamics of the system, especially in materials where ionic conductivity is governed by structural disorder, such as in amorphous solid-state electrolytes. While translational entropy $S_{\text{trans}}$, associated with the motion of ions in space, and vibrational entropy $S_{\text{vib}}$, arising from atomic vibrations, contribute to the total entropy, the configurational entropy becomes the dominant term in systems where ionic transport is heavily influenced by the structural arrangement of ions. The calculated vibrational entropy value is in the range of 1 to $3mk_{\mathrm{B}}$, whereas configurational entropy is in the order of $1k_{\mathrm{B}}$. Therefore, except configurational entropy, other three entropies are negligibly small and impacting on stability.
The Gibbs free energy $\Delta G$ of a system can be written as $\Delta G=H-TS_{\text{total}}$, where $H$ is the enthalpy, $T$ is the temperature, and $S_{\text{total}}$ is the total entropy. Focusing on the configurational entropy contribution, the Gibbs free energy mostly depends on $S_{\text{config}}$. When $S_{\text{config}}$ increases, $\Delta G_{\text{config}}$ decreases, which indicates a more thermodynamically favorable state for ionic transport.
@@
The configurational entropy as a function of $r_{\mathrm{cut}}$ is plotted in Fig. 7 that shows Ti doping at 10% exhibited the highest entropy, followed by 20%. The increased entropy in these cases correlates with higher structural disorder with more stable four coordinated Li atoms (Section 3.5.1 and Fig. 6), which enhances the free-volume diffusion mechanism. At the same time, higher configurational entropy corresponds to a decrease in Gibbs free energy, making the system more thermodynamically favorable for Li-ion diffusion. Specifically, the Ti doping levels of 10% and 20% reduces the Gibbs free energy and they enhance the driving force for ionic transport. This demonstrates the importance of Ti doping concentration in tuning the Li-ion transport mechanism.

### v2

# Mechanisms and Stability of Li Dynamics in Amorphous Li-Ti-P-S-Based Mixed Ionic–Electronic Conductors: A Machine Learning Molecular Dynamics Study

@@ Abstract
Mixed ionic-electronic conductors (MIECs) exhibit both high ionic and electronic conductivity to improve the battery performance. In this work, we investigate the mechanism and stability of transport channels in our recently developed MIEC material, amorphous Ti-doped lithium phosphorus sulfide (LPS), using molecular dynamics (MD) simulations with a 99% accurate machine-learning force field (MLFF) trained on ab-initio MD data. The achieved MLFF helps efficient large-scale MD simulations on LPS with three Ti concentrations (10%, 20%, and 30%) and six temperatures (25oC to 225oC) to calculate ionic conductivity, activation energy, Li-ion transport mechanism, and configurational entropy. Results show that ionic conductivities and activation energies are consistent with our recent experimental values. Moreover, Li-ion transport occurs via free-volume diffusion facilitated by the formation of disordered Li-S polyhedra. The enhanced stability of transport channels at 10% and 20% Ti doping, compared to 0% and 30%, is observed by analyzing the vibrational and configurational entropy of these disordered Li-S polyhedra. Overall, this study highlights the utility of MLFF-based large-scale MD simulations in explaining the transport mechanism and the stability of Li-ion in Ti-doped LPS electrolyte with significant computational efficiency.
@@ 1 Introduction
A MIEC in crystalline or amorphous state represents a promising class of materials for next-generation energy storage devices, including lithium-ion and lithium-sulfur (Li-S) batteries for more than one decade[7, 9, 10]. These materials exhibit both ionic and electronic conductivity, enabling them to play a dual role in facilitating lithium-ion transport while simultaneously providing pathways for electronic charge transfer. Unlike traditional solid electrolytes, which are purely ionic conductors, MIECs improve overall cell efficiency by addressing critical challenges such as sluggish redox kinetics, insulating discharge products, and interfacial instabilities[11]. Especially, the amorphous structure further enhances their ionic conductivity by reducing crystallographic barriers to ion migration, making them highly attractive for solid-state applications[12].
@@
Our approach in this work involves the integration of DFT-based $ab-initio$ molecular dynamics (AIMD), machine learning force filed based on deep neural networks, and classical molecular dynamics to investigate the structural and Li-ion dynamics in Li-Ti-P-S electrolyte. This integrated approach is collectively referred to as deep learning molecular dynamics (DLMD) simulations and its force field is called deep learning potential (DLP) [17, 18]. With this, we organize the manuscript as follows: Section 2 describes detailed simulation techniques, as illustrated in Fig. 1. Section 3 is composed of results and discussions of mean square displacement (MSD), diffusion coefficient, ionic conductivity, activation energy, transport mechanism, and stability of Li-ion ions. Finally, we conclude our results and discussions in Section 4.
Figure 1: (a) Schematic diagram of the protocol of DLMD development (see Section 2.3), where R, F, E, and DL represent atomic coordinates, forces, energies, and deep learning, respectively. (b-e) Ball and stick models of amorphous LPS(b), LPS:Ti10%(c), LPS:Ti20%(d), and LPS:Ti30%(e), respectively. (f-j) Accuracy of trained force field model with DFT data versus predicted data such as energy(f) and forces along Fx, Fy, and Fz of Li(g), P(h), S(i), and Ti(j).
@@ 2.3 DLMD simulation
The training process consisted of $2\times 10^{4}$ steps, utilizing mean squared error for both energies and forces as the loss function at each training step. The optimization was facilitated by the Adam optimizer, initiating with a learning rate of $10^{-3}$ and concluding at $10^{-8}$, with a decay parameter set to 5000. With these specified parameters, a dataset comprising more than 2 million training samples and 20000 test samples from diverse trajectories was employed to construct the DLP model. The calculated loss function parameters such as mean absolute error and root mean square error during validation of the developed DLP model are tabulated in Table-1. Accuracy of the predicted data of force and energy is more than 99 % and data points of predicted versus trained are shown in Fig. 1f to Fig. 1j.
@@ 3.1 Structure optimization
Using the developed DLP, DLMD simulations were performed in the isothermal-isobaric (NPT) ensemble to control the temperature and pressure of the system and optimize a large structure consisting of 12,000 atoms, designed from the output of AIMD. The simulation process began with an initial phase at a low temperature of 1 K for 2 ps to stabilize the initial configuration. Subsequently, the system was gradually heated from 1 K to 1000 K over another 2 ps, allowing a smooth transition to a higher thermal state. After reaching 1000 K, the system was equilibrated at this elevated temperature for 3 ps to ensure thermal and pressure stability. A controlled cooling phase was then implemented, reducing the temperature from 1000 K to 300 K over 3 ps. Finally, the system was equilibrated at 300 K for 50 ps to simulate standard room-temperature conditions. The temperature, potential energy, pressure, and volume profiles with respect to time for LPS:Ti00%,10%,20%,30% systems are plotted in Fig. 2 (a-d) and their optimized structures are shown in Fig. 2(e-h). All their simulations were done using the LAMMPS simulation package [28, 29, 30] coupled with the DeepMD plugin [31].
With the optimized large structure, DLMD simulations were conducted in the NVT ensemble to investigate the Li-dynamic properties. Simulations were performed at eight different temperatures such as 300, 320, 340, 360, 380, 400, 450, and 500 K (25oC to 225oC), controlled using the Nosé-Hoover thermostat. The temperature damping parameter was set to 100 fs, and a uniform integration time step of 1 fs was employed for all simulations, extending over a total duration of 3.0 ns. The small and large cells from the AIMD and DLMD simulations were visualized using VESTA software[32], as shown in Fig. 1 (b-e) and Fig. 2 (e-h), respectively. Additionally, the trajectories of the AIMD and DLMD simulations were analyzed and visualized using OVITO software.[33]
@@ 3.3 Ionic Conductivity and Activation Energy
Figure 5: Ball and stick model for moving Li-ion(Red, Li*) with respect to time step of LPS:Ti00%(a), LPS:Ti10%(b), LPS:Ti20%(c), and LPS:Ti30%(d). Here, the cubic boundaries of the ball and stick models are psudo-boundaries that are only to show the movement of one particular Li-ion in a large cell. The Li-ion movement in 2D plane ($xy$ and $xz$ in Å) with respect to time steps is plotted in (e-l).
@@ 3.4 Li-ion Transport Mechanism
Furthermore, the free-volume diffusion process in LPS involves the formation of disordered Li-Sn ($n$ = 1 to 6) polyhedra within the electrolyte matrix. The disordered nature of these polyhedra reduces structural constraints which results in variations in the migration speed of Li-ions and affects the stability of Li-ion migration channels. Therefore, a quantitative analysis of Li-S polyhedra in Ti-doped LPS is crucial for understanding the stability and effectiveness of Li-ion transport channels.
@@ 3.5.2 Configurational Entropy
where $S_{\text{trans}}$ is the translational entropy, associated with the motion of ions in space; $S_{\text{vib}}$ is the vibrational entropy, arising from atomic vibrations; $S_{\text{elec}}$ is the electronic entropy, accounting for electronic excitations; and $S_{\text{config}}$ is the configurational entropy, which in this study explicitly refers to the degree of disorder in the Li–S polyhedra within the amorphous Li-Ti-P-S matrix.
The vibrational and configurational entropy in Equation 8 was derived from the probabilities of different coordination environments using the Boltzmann entropy formula:
@@
The calculated vibrational entropy value is in the range of 1 to $3mk_{\mathrm{B}}$, whereas configurational entropy is in the order of $1k_{\mathrm{B}}$. This reveals that the configurational entropy becomes particularly significant in systems where ionic transport is strongly influenced by the structural arrangement and disorder of these polyhedra, whereas the contributions from vibrational, translational, and electronic entropies are negligibly small. The Gibbs free energy $\Delta G$ of a system can be written as $\Delta G=H-TS_{\text{total}}$, where $H$ is the enthalpy, $T$ is the temperature, and $S_{\text{total}}$ is the total entropy. Focusing on the configurational entropy contribution, the Gibbs free energy mostly depends on $S_{\text{config}}$. When $S_{\text{config}}$ increases, $\Delta G_{\text{config}}$ decreases, which indicates a more thermodynamically favorable state for ionic transport.
@@
The configurational entropy as a function of $r_{\mathrm{cut}}$ is plotted in Fig. 7 which shows Ti doping at 10% exhibited the highest entropy, followed by 20%. The increased entropy in these cases correlates with higher structural disorder with more stable four-coordinated Li atoms (Section 3.5.1 and Fig. 6), which enhances the free-volume diffusion mechanism. At the same time, higher configurational entropy corresponds to a decrease in Gibbs free energy, making the system more thermodynamically favorable for Li-ion diffusion. Specifically, 10% and 20% of Ti doping reduce the Gibbs free energy and they enhance the driving force for ionic transport. This demonstrates the importance of Ti doping concentration in tuning the Li-ion transport mechanism.